OpenStudy (anonymous):

Explain, in complete sentences, how you would use the graphing method to solve the following system of inequalities and what the final solution will look like on the coordinate plane. x – 2y –8 7x + y < –6

OpenStudy (anonymous):

@johnweldon1993 can you help with this one also? :)

OpenStudy (johnweldon1993):

Missing the inequality sign on the 1st equation...

OpenStudy (anonymous):

x-2y(greaterthan or =to) -8

OpenStudy (johnweldon1993):

Okay...so first step would be...solve each equation for 'y'....what equations do you get when you do that?

OpenStudy (anonymous):

y(greaterthanoreaqualto)4-x y(lessthan)-6-7x

OpenStudy (johnweldon1993):

\[y < -7x - 6\] is correct But the other one...we'll do together \[x - 2y \ge -8\] subtract x from both sides \[-2y \ge -x - 8\] divide everything by -2 ***remember when dividing by a negative number to switch the direction of the sign*** \[y \le \frac{ 1 }{ 2 }x + 4\] understand that part?

OpenStudy (anonymous):

yess

OpenStudy (johnweldon1993):

Okay...so now when graphing....use the fact that \[y < -7x - 6\] -6 is the y-intercept and -7 is the slope and \[y \le \frac{ 1 }{ 2 }x + 4\] 4 is the y-intercept and 1/2 is the slope

OpenStudy (anonymous):

okay and thats it?

OpenStudy (johnweldon1993):

Almost....you graph those lines....then you need to figure out where the solutions lie.... notice the *less than* and *less than or equal to* signs This means you would shade BELOW the lines...and where the 2 shades overlap...are your possible answers

OpenStudy (johnweldon1993):

I'll give you an example for this problem...

OpenStudy (johnweldon1993):